OpenStudy (anonymous):

Covert 189 degre 51' to degree and decimal fractions

OpenStudy (jdoe0001):

there are 60 minutes in 1 degree so how many degrees do you think are there in 60 minutes?

OpenStudy (jdoe0001):

in 60 mins are 1 degree or 60/60 = 1 so in 120 mins there are 120/60 = 2 in 180 mins there are 180/60 = 3 so how many do you think there are in 51 mins?

OpenStudy (anonymous):

\[189+\frac{51}{60}=\frac{3797}{20}=189.85 \]